# Hidden monotonicity and canonical transformations for mean field games and master equations

### v1

@@ 2.2.1. Our results and previous results on the master equation involving displacement semi-monotone data
We note that Theorem 1.1 shows that we have a global well-posedness theory for the master equation as long as $G$ is displacement semi-monotone and the corresponding $\tilde{H}$ is displacement monotone. In particular, it is enough for these to satisfy the ‘first order’ monotonicity conditions, in the sense of Definition 2.1(1) and (2.1). Therefore, Theorem 1.1 together with the well-posedness results from [BMM23] provide a more general result than the one in [MZ22, Theorem 5.6].
2.2.2. Our results and previous results on the master equation involving anti-monotone data
@@
$\int_{\mathbb{R}^{d}}\left\{\left|\partial_{xx}G(x,\mu)\xi(x)+\frac{\lambda_{0}}{2}\xi(x)\right|^{2}+\lambda_{2}\left|\int_{\mathbb{R}^{d}}\partial_{x\mu}G(x,\mu,\tilde{x})\xi(\tilde{x}){\mathrm{d}}\mu(\tilde{x})+\frac{\lambda_{1}}{2\lambda_{2}}\xi(x)\right|^{2}\right\}{\mathrm{d}}\mu(x)$
$\leq\left(\lambda_{3}+\left(\frac{\lambda_{0}}{2}\right)^{2}+\lambda_{2}\left(\frac{\lambda_{1}}{2\lambda_{2}}\right)^{2}\right)\int_{\mathbb{R}^{d}}\left|\xi(x)\right|^{2}{\mathrm{d}}\mu(x)$
@@
$\left|\int_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\langle\partial_{x\mu}G(x,\mu,\tilde{x})\xi(x),\xi(\tilde{x})\rangle{\mathrm{d}}\mu(x){\mathrm{d}}\mu(\tilde{x})\right|\leq\left(\frac{\left|\lambda_{1}\right|}{2\lambda_{2}}+\sqrt{\frac{\lambda_{3}}{\lambda_{2}}+\frac{{\lambda_{0}}^{2}}{4\lambda_{2}}+{\left(\frac{\lambda_{1}}{2\lambda_{2}}\right)^{2}}}\right)\int_{\mathbb{R}^{d}}\left|\xi(x)\right|^{2}{\mathrm{d}}\mu(x)$
@@
$\left|\int_{\mathbb{R}^{d}}\langle\partial_{xx}G(x,\mu)\xi(x),\xi(x)\rangle{\mathrm{d}}\mu(x)\right|\leq\left(\frac{\left|\lambda_{0}\right|}{2}+\sqrt{\lambda_{3}+\left(\frac{\lambda_{0}}{2}\right)^{2}+\lambda_{2}\left(\frac{\lambda_{1}}{2\lambda_{2}}\right)^{2}}\right)\int_{\mathbb{R}^{d}}{\mathrm{d}}\mu(x)\left|\xi(x)\right|^{2}$
@@
With the choice of $v:=\frac{\lambda_{1}}{2\lambda_{2}}\xi(x)$ and $w:=\int_{\mathbb{R}^{d}}\partial_{x\mu}G(x,\mu,\tilde{x})\xi(\tilde{x}){\mathrm{d}}\mu(\tilde{x})$, we obtain
@@
$\leq\int_{\mathbb{R}^{d}}\left\{\left(\frac{C}{2}+1\right)\left|\frac{\lambda_{1}}{2\lambda_{2}}\xi(x)\right|^{2}+\frac{1}{2C}\left|\int_{\mathbb{R}^{d}}\partial_{x\mu}G(x,\mu,\tilde{x})\xi(\tilde{x}){\mathrm{d}}\mu(\tilde{x})+\frac{\lambda_{1}}{2\lambda_{2}}\xi(x)\right|^{2}\right\}{\mathrm{d}}\mu(x)$
$=\int_{\mathbb{R}^{d}}\left\{\left(\frac{C}{2}+1\right)\left|\frac{\lambda_{1}}{2\lambda_{2}}\xi(x)\right|^{2}+\frac{1}{2C\lambda_{2}}\left(\lambda_{2}\left|\int_{\mathbb{R}^{d}}\partial_{x\mu}G(x,\mu,\tilde{x})\xi(\tilde{x}){\mathrm{d}}\mu(\tilde{x})+\frac{\lambda_{1}}{2\lambda_{2}}\xi(x)\right|^{2}\right)\right\}{\mathrm{d}}\mu(x)$
$\leq\int_{\mathbb{R}^{d}}\left\{\left(\frac{C}{2}+1\right)\left|\frac{\lambda_{1}}{2\lambda_{2}}\xi(x)\right|^{2}+\frac{1}{2C\lambda_{2}}\left(\lambda_{3}+\left(\frac{\lambda_{0}}{2}\right)^{2}+\lambda_{2}\left(\frac{\lambda_{1}}{2\lambda_{2}}\right)^{2}\right)\left|\xi(x)\right|^{2}\right\}{\mathrm{d}}\mu(x)$
@@
$f(\lambda):=\frac{5{\left|\lambda_{1}\right|}}{4\lambda_{2}}+1+{\frac{\lambda_{3}}{2\lambda_{2}}}+\frac{\lambda_{0}}{4\lambda_{2}}+\frac{5\lambda_{0}}{4}+\frac{\lambda_{3}}{2}+\frac{{\left|\lambda_{1}\right|}}{4}=1+\frac{1}{2}\left(\frac{5\lambda_{0}}{2}+\frac{{\left|\lambda_{1}\right|}}{2}+\lambda_{3}\right)+\frac{1}{2\lambda_{2}}\left(\frac{\lambda_{0}}{2}+\frac{5{\left|\lambda_{1}\right|}}{2}+\lambda_{3}\right),$
@@
$\underline{\kappa}({\partial_{xp}H})$ $\geq\underline{\kappa}(A_{0})-\left|\partial_{xp}H_{0}\right|\geq 3L_{2}^{H_{0}}+\frac{1}{2}\left|\partial_{p\mu}H\right|-\left|\partial_{xp}H_{0}\right|+\frac{\sqrt{K_{H}}}{2}L_{2}^{H_{0}}+\sqrt{\left|\partial_{pp}H\right|\left|\partial_{xx}H_{0}\right|}$
@@
(${}^{\prime}$) $\underline{\kappa}(A_{0})\geq(1+\underline{\kappa}(A_{1}^{-1}A_{2}))L_{2}^{H_{0}},$
although they probably meant to assume (2.7)111 The $\underline{\kappa}$ on the right-hand side is likely a typo as in the fourth to last line on [MZ24, page 15] the authors need to use $\bar{\kappa}(A_{1}^{-1}A_{2})$. Furthermore we see $\bar{\kappa}$ appearing correctly also in a similar assumption, [MZ22, (6.3)]. .
@@
Furthermore we see from the second inequality in [MZ24, (7.2)] that $\bar{\gamma}\underline{\kappa}(A_{0})\geq\left|\partial_{xx}H\right|$. By the assumption (i) of [MZ24, Theorem 7.1] we have that $\bar{\gamma}$ satisfies [MZ24, (4.2)] in which the first inequality implies that $\lambda_{0}>\frac{\bar{\gamma}^{2}}{4\underline{\gamma}}-\frac{8\lambda_{3}}{4\underline{\gamma}}$. Hence we obtain $(4\underline{\gamma}\lambda_{0}+8\lambda_{3})\geq\bar{\gamma}^{2}$. It is clear that $2f(\lambda)\geq\lambda_{0}$ and $2f(\lambda)\geq\lambda_{3}$, therefore we get $16f(\lambda)(1+\underline{\gamma})\geq\bar{\gamma}^{2}$. Since $\underline{\gamma}<\bar{\gamma}$ by assumption (i) of [MZ24, Theorem 7.1] and $1<\bar{\gamma}$ by the same assumption we get $2\bar{\gamma}\geq 1+\underline{\gamma}$ and so we obtain $32f(\lambda)\geq\bar{\gamma}$. Hence we get

### v2

@@ 3.1. Our results and previous results on the master equation involving displacement semi-monotone data
We note that Theorem 3.2 shows that we have a global well-posedness theory for the master equation as long as $G$ is displacement semi-monotone and the corresponding $\tilde{H}$ is displacement monotone. In particular, it is enough for these to satisfy the ‘first order’ monotonicity conditions, in the sense of Definition 2.1(1) and (2.1). Therefore, Theorem 3.2 together with the well-posedness results from [BMM25] provide a more general result than the one in [MZ22, Theorem 5.6].
3.2. Our results and previous results on the master equation involving anti-monotone data
@@
$\int_{\mathbb{R}^{d}}\bigg\{\left|\partial_{xx}G(x,\mu)\xi(x)+\frac{\lambda_{0}}{2}\xi(x)\right|^{2}$
$+\lambda_{2}\left|\int_{\mathbb{R}^{d}}\partial_{x\mu}G(x,\mu,\tilde{x})\xi(\tilde{x}){\mathrm{d}}\mu(\tilde{x})+\frac{\lambda_{1}}{2\lambda_{2}}\xi(x)\right|^{2}\bigg\}{\mathrm{d}}\mu(x)$
$\leq\left(\lambda_{3}+\left(\frac{\lambda_{0}}{2}\right)^{2}+\lambda_{2}\left(\frac{\lambda_{1}}{2\lambda_{2}}\right)^{2}\right)\int_{\mathbb{R}^{d}}\left|\xi(x)\right|^{2}{\mathrm{d}}\mu(x).$
@@
[…]
$\qquad\qquad\leq\left(\frac{\left|\lambda_{1}\right|}{2\lambda_{2}}+\sqrt{\frac{\lambda_{3}}{\lambda_{2}}+\frac{{\lambda_{0}}^{2}}{4\lambda_{2}}+{\left(\frac{\lambda_{1}}{2\lambda_{2}}\right)^{2}}}\right)\int_{\mathbb{R}^{d}}\left|\xi(x)\right|^{2}{\mathrm{d}}\mu(x)$
@@
$\left|\int_{\mathbb{R}^{d}}\langle\partial_{xx}G(x,\mu)\xi(x),\xi(x)\rangle{\mathrm{d}}\mu(x)\right|$
$\qquad\qquad\leq\left(\frac{\left|\lambda_{0}\right|}{2}+\sqrt{\lambda_{3}+\left(\frac{\lambda_{0}}{2}\right)^{2}+\lambda_{2}\left(\frac{\lambda_{1}}{2\lambda_{2}}\right)^{2}}\right)\int_{\mathbb{R}^{d}}{\mathrm{d}}\mu(x)\left|\xi(x)\right|^{2}.$
@@
[…]
@@
$\leq\int_{\mathbb{R}^{d}}\bigg\{\left(\frac{C}{2}+1\right)\left|\frac{\lambda_{1}}{2\lambda_{2}}\xi(x)\right|^{2}$
$+\frac{1}{2C}\left|\int_{\mathbb{R}^{d}}\partial_{x\mu}G(x,\mu,\tilde{x})\xi(\tilde{x}){\mathrm{d}}\mu(\tilde{x})+\frac{\lambda_{1}}{2\lambda_{2}}\xi(x)\right|^{2}\bigg\}{\mathrm{d}}\mu(x)$
$=\int_{\mathbb{R}^{d}}\bigg\{\left(\frac{C}{2}+1\right)\left|\frac{\lambda_{1}}{2\lambda_{2}}\xi(x)\right|^{2}$
$+\frac{1}{2C\lambda_{2}}\left(\lambda_{2}\left|\int_{\mathbb{R}^{d}}\partial_{x\mu}G(x,\mu,\tilde{x})\xi(\tilde{x}){\mathrm{d}}\mu(\tilde{x})+\frac{\lambda_{1}}{2\lambda_{2}}\xi(x)\right|^{2}\right)\bigg\}{\mathrm{d}}\mu(x)$
$\leq\int_{\mathbb{R}^{d}}\bigg\{\left(\frac{C}{2}+1\right)\left|\frac{\lambda_{1}}{2\lambda_{2}}\xi(x)\right|^{2}$
$+\frac{1}{2C\lambda_{2}}\left(\lambda_{3}+\left(\frac{\lambda_{0}}{2}\right)^{2}+\lambda_{2}\left(\frac{\lambda_{1}}{2\lambda_{2}}\right)^{2}\right)\left|\xi(x)\right|^{2}\bigg\}{\mathrm{d}}\mu(x)$
@@
$f(\lambda)$ $:=\frac{5{\left|\lambda_{1}\right|}}{4\lambda_{2}}+1+{\frac{\lambda_{3}}{2\lambda_{2}}}+\frac{\lambda_{0}}{4\lambda_{2}}+\frac{5\lambda_{0}}{4}+\frac{\lambda_{3}}{2}+\frac{{\left|\lambda_{1}\right|}}{4}$
$=1+\frac{1}{2}\left(\frac{5\lambda_{0}}{2}+\frac{{\left|\lambda_{1}\right|}}{2}+\lambda_{3}\right)+\frac{1}{2\lambda_{2}}\left(\frac{\lambda_{0}}{2}+\frac{5{\left|\lambda_{1}\right|}}{2}+\lambda_{3}\right),$
@@
$\underline{\kappa}({\partial_{xp}H})$ $\geq\underline{\kappa}(A_{0})-\left|\partial_{xp}H_{0}\right|\geq 3L_{2}^{H_{0}}+\frac{1}{2}\left|\partial_{p\mu}H\right|-\left|\partial_{xp}H_{0}\right|+\frac{\sqrt{K_{H}}}{2}L_{2}^{H_{0}}$
$+\sqrt{\left|\partial_{pp}H\right|\left|\partial_{xx}H_{0}\right|}$
@@
(3.5) $\underline{\kappa}(A_{0})\geq(1+\underline{\kappa}(A_{1}^{-1}A_{2}))L_{2}^{H_{0}},$
although they probably meant to assume (3.5) [^1] The $\underline{\kappa}$ on the right-hand side is likely a typo as in the fourth to last line on [MZ24, page 15] the authors need to use $\bar{\kappa}(A_{1}^{-1}A_{2})$. Furthermore we see $\bar{\kappa}$ appearing correctly also in a similar assumption, [MZ22, (6.3)]. .
@@
Furthermore we see from the second inequality in [MZ24, (7.2)] that
$\bar{\gamma}\underline{\kappa}(A_{0})\geq\left|\partial_{xx}H\right|.$
By the assumption (i) of [MZ24, Theorem 7.1] we have that $\bar{\gamma}$ satisfies [MZ24, (4.2)] in which the first inequality implies that $\lambda_{0}>\frac{\bar{\gamma}^{2}}{4\underline{\gamma}}-\frac{8\lambda_{3}}{4\underline{\gamma}}$. Hence we obtain $(4\underline{\gamma}\lambda_{0}+8\lambda_{3})\geq\bar{\gamma}^{2}$. It is clear that $2f(\lambda)\geq\lambda_{0}$ and $2f(\lambda)\geq\lambda_{3}$, therefore we get $16f(\lambda)(1+\underline{\gamma})\geq\bar{\gamma}^{2}$. Since $\underline{\gamma}<\bar{\gamma}$ by assumption (i) of [MZ24, Theorem 7.1] and $1<\bar{\gamma}$ by the same assumption we get $2\bar{\gamma}\geq 1+\underline{\gamma}$ and so we obtain $32f(\lambda)\geq\bar{\gamma}$. Hence we get
@@
Data availability statement. No data was generated for the purposes of this research.
Declarations. The authors declare that there are no conflicts or competing interests.